Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b
and
mx + y = 22

Answer by venugopalramana(3286)

Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b..PUT (-3,4) IN THIS
5*-3+7*4=B
B=-15+28=13
mx + y = 22...PUT (-3,4) IN THIS
M*-3+4=22
-3M=22-4=18
M=-18/3=-6